A. C. Neilsen reported that children between the ages of 2 and 5 watch an average of 25 hours of television per week. Assume the variable is normally distributed and the standard deviation is 3 hours. If 20 children between the ages of 2 and 5 are randomly selected, find . Give answer to three decimal places.
Answer #1

The following information about the mean and standard deviation has been provided:

μ=25, σ=3, n=20

We need to compute Pr(Xˉ≥26.3).

The corresponding z-value needed to be computed is:

Z = \frac{\bar X-\mu}{\sigma/\sqrt{n}} = \frac{ 26.3-25}{ 3/\sqrt{ 20}} = 1.9379

Therefore, we get that

\Pr(\bar X \geq 26.3) = \Pr\left(Z \ge \frac{ 26.3-25}{ 3/\sqrt{ 20}}\right) = \Pr(Z \ge 1.9379)

=1-0.9737=0.026

Please like